Chell:

A train is moving 4.82 m/s when it begins to accelerate. After 24.8s it is moving 7.09m/s how far did it move in that time

SemiDefinite:

sounds like SUVAT, again: - u = 4.82 m/s - t = 24.8s - v = 7.09m/s - s = ??